# The 1/3 Geometric Constant: Scale Invariance and the Origin of ’Missing Energy’ in 3D Quantum Fragmentation

### v2

@@ Abstract
We report a fundamental geometric constraint on the detection of kinetic energy release (KER) in three-dimensional quantum fragmentation. By investigating the sudden dissociation of Slater-type orbitals—a physically motivated basis for localized states—we demonstrate that the ratio of the peak detected energy to the integrated mean, $R_{E}=E_{\text{peak}}/\langle E\rangle$, is strictly bounded below $0.5$ for all physical parameter spaces. We systematically map the behavior of $R_{E}$ across the orbital localization ($\zeta$) and effective repulsive charge ($Q$) landscape. Our results show that while $R_{E}$ exhibits sensitivity to the force-field geometry at atomic scales, it converges to a remarkably stable, scale-invariant constant of $\approx 0.33$ in the high-localization limit ($M\propto\zeta$) characteristic of subatomic ground states ($n=1,Q=1$) . This $1/3$ geometric constant provides a provocative first-principles interpretation of the historical “missing energy” problem in nuclear physics. We suggest that the $1/3$ average energy ratio observed in Beta decay spectra may be a topological artifact of the $4\pi r^{2}$ volume weighting inherent to 3D wavefunctions, rather than exclusively a signature of undetected mass. This work establishes $R_{E}<0.5$ as a universal geometric baseline for quantum simulations and offers a new framework for reconstructing the true energy budget in subatomic calorimetry by correcting for the intrinsic masking effects of 3D radial geometry.
@@ I Introduction
The sudden approximation remains a cornerstone of quantum dynamics, providing a robust framework for systems where an instantaneous change in the Hamiltonian occurs before the wavefunction can spatially redistribute Makarov (2016); Eichler and Meyerhof (1995). This regime is particularly critical in the study of Coulomb explosions and nuclear dissociations, where rapid transitions from attractive to repulsive potentials lead to the fragmentation of the system. Historically, classical and 1D semiclassical models have been the primary tools for interpreting the resulting kinetic energy release (KER) Yue and Madsen (2013). However, these models frequently struggle to replicate the precise energy distributions observed in high-resolution experiments, often overestimating the detected peak energy and failing to account for significant portions of the integrated energy budget.
In this work, we demonstrate that the persistent discrepancy between theoretical energy conservation and experimental observation is a fundamental consequence of 3D quantum geometry. By utilizing a local energy framework applied to Slater-type orbitals (STOs)—the physical standard for localized quantum systems—we show that the energy distribution $P(E)$ is governed by the $4\pi r^{2}$ radial volume weighting. A primary finding of this study is the “0.5 Rule”: the ratio of the peak detected energy to the integrated statistical mean ($R_{E}=E_{\text{peak}}/\langle E\rangle$) is strictly bounded below $0.5$ for ground-state configurations. This shift occurs because the 3D volume element suppresses the probability density near the potential energy singularity at the origin, a geometric effect that 1D simulations inherently fail to capture.
To establish the universality of this phenomenon, we perform an exhaustive parametric study across the principal quantum number ($n$), the orbital exponent ($\zeta$), and the repulsive charge ($Q$). Crucially, we investigate the scaling limit where the particle mass $M$ is proportional to the localization parameter $\zeta$. Under these conditions, we uncover a scale-invariant regime where the ratio $R_{E}$ converges to a stable constant, which is $\approx 0.33$ for a repulsive unit charge ($Q=1$).
This $1/3$ geometric constant provides a provocative re-interpretation of the “missing energy” problem in subatomic physics. Just as the beta-decay crisis of the early 20th century necessitated the neutrino to satisfy energy conservation based on an observed $1/3$ average energy ratio Pauli (1930); Ellis and Wooster (1927), our results suggest that such imbalances may be an intrinsic signature of the initial state’s 3D geometry. If the detected peak inherently captures only one-third of the true interaction energy due to radial sequestering, then “missing” energy may represent the recovered high-momentum components of the primary fragment rather than undetected mass. This work establishes a necessary geometric baseline for interpreting energy calorimetry across atomic, nuclear, and high-energy scales.
II Theoretical Methodology
II.1 The Sudden Approximation and Potential Inversion
@@
II.2 Bohmian Local Energy Formulation
To determine the kinetic energy distribution of the resulting fragments, we utilize the de Broglie-Bohm formulation of quantum mechanics. In this framework, the total energy $E(\mathbf{r})$ of a particle at a specific spatial coordinate $\mathbf{r}$ is defined by the sum of the repulsive potential and the local kinetic energy density $T_{local}(\mathbf{r})$:
@@ II.2 Bohmian Local Energy Formulation
$T_{local}(\mathbf{r})=-\frac{\hbar^{2}}{2m}\frac{\nabla^{2}\psi_{0}(\mathbf{r})}{\psi_{0}(\mathbf{r})}$ (4)
@@
$T_{local}(r)=\frac{\hbar^{2}}{2m}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)$ (5)
@@
$E(r)=\frac{\hbar^{2}}{2m}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)+\frac{Q}{r}.$ (6)
II.3 Numerical Implementation and Statistical Weighting
@@
III Distribution property
Figure 1: Comparison of normalized energy distributions P(E) between the present model (solid) and tRecX TDSE simulations (dashed). Results are shown for repulsion charges $Q=0,0.5$, and using an initial 1s hydrogenic orbital ($\zeta=1,l=0$). Peak positions show high consistency across all charge values.
To validate the predictive power of our 3D Bohmian-based model, we compared the resulting energy distributions P(E) against full quantum simulations performed via the tRecX Scrinzi (2022); Zhu and Scrinzi (2020); Zhu (2021, 2020); Majety and Scrinzi (2015); Zielinski et al. (2016) software package. We utilized a standard 1s Slater orbital ($\zeta=1$,$l=0$) as the initial state and varied the repulsive Coulomb charge $Q$.
As shown in Figure 1, the peak positions calculated by our method are in excellent agreement with the tRecX benchmarks across all charge values ($Q\in\{0,0.5,1\}$), the shape of our distribution is also similar to tRecX computation. The deviations in the distribution tails are attributed to the numerical constraints of grid-based solvers. Specifically, the truncation of the Coulomb potential and the use of polynomial quadrature grids in tRecX often struggle to resolve the singularity at $r\to 0$, whereas our analytical local-energy approach maintains precision in the near-nucleus region. This consistency in peak prediction confirms that our simplified method effectively captures the primary experimental signatures of the Coulomb explosion.
Figure 2: Parametric evolution of the energy distribution $P(E)$. (Upper) Variation with repulsive Coulomb charge $Q\in[0,2]$ for a fixed 1s orbital ($\zeta=1$,$l=0$). Both the peak position and spectral width increase with Q. (Lower) Variation with orbital exponent $\zeta\in[1,4]$ at a fixed charge $Q=1$. The distribution broadens more rapidly with increasing $\zeta$ compared to $Q$, highlighting the sensitivity of the KER to initial state localization.
@@ III Distribution property
As illustrated in Figure 2 (Upper), increasing the repulsive charge Q leads to a systematic shift in the energy distribution $P(E)$ toward higher energies. This is accompanied by a significant broadening of the spectral width. Since the local energy is defined as $E(r)=T_{local}(r)+Q/r$, a larger $Q$ increases the potential gradient across the spatial extent of the orbital, thereby spreading the resulting energy values over a wider range.
@@
In Figure 2 (Lower), we examine the effect of the orbital exponent $\zeta$ while maintaining a constant charge ($Q=1$). We observe that the peak position of $P(E)$ is highly sensitive to the spatial localization of the initial state. As increases, the wavefunction becomes more tightly bound, which simultaneously increases the local kinetic energy density $T_{local}\propto\zeta^{2}$ and places the particle density in regions of higher potential energy. Interestingly, the distribution broadens much more rapidly with respect to $\zeta$ than it does with respect to Q. This suggests that the internal ”quantum pressure” (local kinetic energy) of the initial state plays a more dominant role in determining the final energy dispersion than the external repulsive force itself.
To identify a universal analytical form for the fragment energy spectrum, we performed a regression analysis on the energy distribution P(E) using several statistical models. Figure 3 illustrates the results for a ground-state hydrogenic orbital ($\zeta=1,Q=1$) and a more localized, higher-charge configuration ($\zeta=5.2,Q=1.6$).
We compared the computed data (blue) against three candidate distributions: a radial-type distribution (r-fit), a Gamma distribution, and a χ distribution. As shown in the figure, the r-fit (green) and the Gamma distribution (cyan) produce the highest fidelity, with a residual error of less than 2%. The near-perfect overlap between these two functions confirms that the energy distribution essentially obeys a radial distribution law.
A critical quantitative observation is found in the integrated energy values shown in the plot headers. In both cases, the peak energy $E_{peak}$ is significantly smaller than the total scattering energy Escatter (the integrated average $\langle E\rangle$). Specifically, for the $\zeta=1,Q=1$ case, the ”energy conservation” value $\langle E\rangle$ is driven higher by the potential energy density near the origin, whereas the experimental peak corresponds to the most probable radial position. This discrepancy validates our hypothesis that fragment detectors fundamentally probe the radial geometry of the initial state rather than the integrated energy mean.
@@
Figure 4: Universal scaling of the peak-to-mean energy ratio $R_{E}=E_{peak}/\langle E\rangle$. (Upper) Variance with repulsive charge Q at $\zeta=1$, showing monotonic convergence toward 0.5 as the repulsive potential dominates. (Middle) Variance with orbital exponent $\zeta$ at Q=1, showing a precipitous drop in detection efficiency for localized states. (Lower) Contour map of $R_{E}(\zeta,Q)$. The ratio is strictly bounded below 0.5, with the greatest discrepancy occurring for inner-shell analogs (high $\zeta$). High $\zeta$ values correspond to deeper initial potentials where the ”quantum pressure” ($T_{local}∼\zeta^{2}$) significantly skews the average energy relative to the experimental peak.
IV Scaling Laws and the Energy Detection Discrepancy
@@
IV.1 Influence of Repulsive Charge $Q$
@@
IV.2 Orbital Exponent $\zeta$: Inner-Shell vs. Outer-Shell Dynamics
@@
IV.3 Universal Contour Map and Numerical Stability
@@
IV.4 The Geometric Constant of Detection and Scale Invariance
As illustrated in the upper figure of Figure 5, for a fixed effective charge $Q=1$, the ratio $R_{E}$ remains strikingly constant at $\approx 0.33$ across an order of magnitude of $\zeta$. Given that charge quantization dictates a unit value for single particle interactions, the $R_{E}≈0.33$ limit at $Q=1$ represents the fundamental baseline for 3D fragmentation. Consequently, this $1/3$ constant is often treated as the intrinsic geometric signature of the system, even when the unit charge $Q=1$ is not explicitly specified.This stability suggests that, for a uinit charge $Q=1$, the $1/3$ ratio is not a result of specific dynamical parameters, but is a topological consequence of the $4\pi r^{2}$ volume element in three-dimensional space. In this limit, the ”missing” energy is naturally sequestered in the high-energy tail of the radial distribution.
The broader landscape is presented in the lower figure of Figure 5. The flatness of the $R_{E}$ surface along the $\zeta$-axis confirms that the detection discrepancy is scale-invariant for any interaction strength $Q$. Crucially, even as $Q$ and $\zeta$ vary, the ratio never exceeds the $0.5$ threshold, reinforcing the conclusion that a primary experimental peak can never capture more than half of the total interaction energy in a 3D ground-state system. This convergence to $1/3$ provides a robust geometric baseline that may account for the energy imbalances traditionally attributed to auxiliary particles in nuclear decay.
Figure 5: Scaling behavior of the peak-to-mean ratio $R_{E}$ for the ground-state ($n=1$) fragmentation. (uppere) 2D profile for $Q=1$ across $\zeta\in[0.25,5]$. The ratio converges to a stable plateau at $\approx 0.33$ (indicated by the dashed line), matching the historical $1/3$ average energy ratio. (lower) 3D contour plot of $R_{E}$ as a function of $\zeta$ and $Q$. The results demonstrate that for any given repulsive charge $Q$, the ratio remains invariant with respect to $\zeta$, while strictly obeying the $R_{E}<0.5$ universal geometric bound across the entire parameter space.
IV.5 Empirical Validation: The $H_{2}^{+}$ Energy Discrepancy
The physical validity of the $1/3$ geometric constant is most clearly evidenced by the long-standing energy discrepancy observed in attosecond $H_{2}^{+}$ fragmentation. Classical energy conservation and 1D semiclassical models predict a kinetic energy release (KER) peak at approximately 10101111 eV for the dissociation of $H_{2}^{+}$ from its ground state. However, both high-level TDSE simulations and experimental calorimetry consistently report a primary peak in the 3344 eV range Zhu (2020, 2021).
As shown in our $H_{2}^{+}$ benchmarks, our 3D geometric model replicates this $\approx 7$ eV shift without requiring the inclusion of external field effects or complex multi-photon dynamics. Because the initial $1\sigma_{g}$ state of $H_{2}^{+}$ is highly localized, its radial volume weighting ($4\pi r^{2}$) effectively sequesters the majority of the potential energy in the high-momentum tail of the distribution. Consequently, the most probable detected energy ($E_{\text{peak}}$) represents only one-third of the integrated mean energy ($\langle E\rangle\approx 11$ eV).
This exact alignment between the predicted $0.33$ ratio and the observed $3/11$ ratio in $H_{2}^{+}$ experiments confirms that the ”missing energy” in attosecond molecular physics is a direct manifestation of the 3D radial distribution. Furthermore, our preliminary analysis of excited states ($n>1$) in $H_{2}^{+}$ indicates the emergence of secondary high-energy peaks, a signature of nodal structures in the initial state that further differentiates 3D quantum geometry from 1D approximations.
V Principal Quantum Number Effects and Multi-Peak Phenomena

### v3

@@ Abstract
We report the discovery of a universal geometric constraint on the detection of kinetic energy release (KER) in three-dimensional quantum fragmentation. By investigating the dissociation of localized Slater-type orbitals, we demonstrate that the $4\pi r^{2}$ radial volume element inherently masks the majority of a system’s energy budget, imposing a fundamental peak-to-mean energy bound of $R_{E}<0.5$. We introduce a dimensionless scaling law, $M=\alpha\zeta/Q$, and prove that the resulting energy detection ratio is scale-invariant across twelve orders of magnitude, bridging the gap between atomic and subatomic physics. Remarkably, we identify a “Geometric Zero-Point” at $R_{E}\approx 0.33$, which precisely replicates the 7 eV “missing energy” anomaly observed in attosecond $H_{2}^{+}$ fragmentation benchmarks. Furthermore, we demonstrate that this $1/3$ ratio provides a robust geometric baseline for the historical average-to-endpoint discrepancy in beta decay. Our results suggest that a significant portion of what is historically categorized as “missing energy” may be a topological artifact of 3D quantum geometry rather than an exclusive signature of undetected particles. This work establishes a universal master curve for energy reconstruction and identifies a “detection crisis” in highly localized systems, where the true interaction energy becomes effectively invisible to peak-centric experimental calorimetry.
@@ I Introduction
A central challenge in experimental physics is the reconstruction of the total energy budget from a detected momentum distribution. Across disparate scales—from the electron-volts of molecular dissociation to the mega-electron-volts of nuclear transitions—a persistent anomaly remains: the primary spectral peak often captures only a fraction of the available energy. Historically, this “missing energy” has been the catalyst for some of the most profound discoveries in physics, most notably the hypothesis of the neutrino.
The birth of the neutrino was necessitated by the observation that the average energy of electrons emitted in beta decay was approximately one-third of the total transition energy (the endpoint) Ellis and Wooster (1927). To preserve energy conservation, Pauli famously postulated an invisible particle to carry the balance.
Remarkably, a nearly identical $1/3$ discrepancy has emerged in modern attosecond calorimetry. In the Coulomb fragmentation of $H_{2}^{+}$, classical energy conservation and 1D models predict a kinetic energy release (KER) peak at $\approx 11$ eV Yue and Madsen (2013). However, high-precision experimental measurements and full-dimensional Time-Dependent Schrödinger Equation (TDSE) simulations consistently report a primary peak between 3344 eV Zhu (2020, 2021). This observed ratio of $R_{E}\approx 0.32$ almost perfectly mirrors the historical beta decay discrepancy, despite the absence of neutrinos in molecular fragmentation.
In this work, we demonstrate that these anomalies share a common, purely geometric origin. We prove that for any localized 3D ground-state wavefunction, the $4\pi r^{2}$ volume element inherently suppresses probability near the origin, sequestering the majority of the interaction energy in a high-momentum radial tail. We identify a universal 0.33 geometric landmark that emerges under a specific dimensionless scaling law, $M=\alpha\zeta/Q$, where $M$ is the particle mass, $\zeta$ the orbital localization, and $Q$ the repulsive charge.
Specifically, we show that as a system becomes more massive and more localized—as in the case of nucleons where $\zeta$ increases alongside MM—the detected peak energy remains a small, fixed fraction of the true energy. This allows us to establish a master curve for energy reconstruction, suggesting that much of the energy historically labeled as “missing” is actually a topological artifact of 3D quantum geometry. This framework provides a new tool for particle verification and discovery, shifting the focus from undetected particles to the intrinsic masking effects of the 3D volume element.
II Methods
III Theoretical Methodology
III.1 The Sudden Approximation and Potential Inversion
@@
III.2 Bohmian Local Energy Formulation
To determine the kinetic energy distribution of the resulting fragments, we utilize the de Broglie-Bohm formulation of quantum mechanics. The atomic units $\hbar=m_{e}=e=1$ are applied if not specfied. In this framework, the total energy $E(\mathbf{r})$ of a particle at a specific spatial coordinate $\mathbf{r}$ is defined by the sum of the repulsive potential and the local kinetic energy density $T_{local}(\mathbf{r})$:
@@ III.2 Bohmian Local Energy Formulation
$T_{local}(\mathbf{r})=-\frac{1}{2M}\frac{\nabla^{2}\psi_{0}(\mathbf{r})}{\psi_{0}(\mathbf{r})}$ (4)
@@
$T_{local}(r)=\frac{1}{2M}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)$ (5)
@@
$E(r)=\frac{1}{2M}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)+\frac{Q}{r}.$ (6)
III.3 Numerical Implementation and Statistical Weighting
@@
IV Distribution property
To validate the predictive power of our 3D Bohmian-based model, we compared the resulting energy distributions $P(E)$ against full quantum simulations performed via the tRecX Scrinzi (2022); Zhu and Scrinzi (2020); Zhu (2021, 2020); Majety and Scrinzi (2015); Zielinski et al. (2016) software package. We utilized a standard 1s Slater orbital ($\zeta=1$,$l=0$) as the initial state and varied the repulsive Coulomb charge $Q$. We will atomic units
As shown in upper figure of Fig 1, the peak positions calculated by our method are in excellent agreement with the tRecX benchmarks across all charge values ($Q\in\{0,0.5,1\}$), the shape of our distribution is also similar to tRecX computation. The deviations in the distribution tails are attributed to the numerical constraints of grid-based solvers. Specifically, the truncation of the Coulomb potential and the use of polynomial quadrature grids in tRecX often struggle to resolve the singularity at $r\to 0$, whereas our analytical local-energy approach maintains precision in the near-nucleus region. This consistency in peak prediction confirms that our simplified method effectively captures the primary experimental signatures of the Coulomb explosion.
Figure 1: Comparison of normalized energy distributions P(E) between the present model (solid) and tRecX TDSE simulations (dashed). Results are shown for repulsion charges $Q=0,0.5$, and using an initial 1s hydrogenic orbital ($\zeta=1,l=0$). Peak positions show high consistency across all charge values. Parametric evolution of the energy distribution $P(E)$. (Middle) Variation with repulsive Coulomb charge $Q\in[0,2]$ for a fixed 1s orbital ($\zeta=1$,$l=0$). Both the peak position and spectral width increase with Q. (Lower) Variation with orbital exponent $\zeta\in[1,4]$ at a fixed charge $Q=1$. The distribution broadens more rapidly with increasing $\zeta$ compared to $Q$, highlighting the sensitivity of the KER to initial state localization.
@@ IV Distribution property
As illustrated in Fig. 1 (Upper), increasing the repulsive charge $Q$ leads to a systematic shift in the energy distribution $P(E)$ toward higher energies, where the peak of $P(E)$ is proportional to $Q$. This is accompanied by a broadening to some extend of the spectral width. Since the local energy is defined as $E(r)=T_{local}(r)+Q/r$, a larger $Q$ increases the potential gradient across the spatial extent of the orbital, thereby spreading the resulting energy values over a wider range. The shape of the curve could be fited by a radial distribution, as is illustrated in Sec. A in the appendix, suggesting inherently connection with the radial coordinates.
@@
In Fig. 1 (Lower), we examine the effect of the orbital exponent $\zeta$ while maintaining a constant charge ($Q=1$). We observe that the peak position of $P(E)$ is also proportional to $\zeta$. As $\zeta$ increases, the wavefunction becomes more tightly bound, which simultaneously increases the local kinetic energy density $T_{local}\propto\zeta^{2}$ and places the particle density in regions of higher potential energy. Interestingly, the distribution broadens much more rapidly with respect to $\zeta$ than it does with respect to $Q$. This suggests that the internal ”quantum pressure” (local kinetic energy) of the initial state plays a more dominant role in determining the final energy dispersion than the external repulsive force itself.
@@
Figure 2: Universal scaling of the peak-to-mean energy ratio $R_{E}=E_{\text{peak}}/\langle E\rangle$. (Upper) Variance with repulsive charge $Q$ at $\zeta=1$, showing monotonic convergence toward 0.5 as the repulsive potential dominates. (Middle) Variance with orbital exponent $\zeta$ at $Q=1$, showing a precipitous drop in detection efficiency for localized states. (Lower) Contour map of $R_{E}(\zeta,Q)$. The ratio is strictly bounded below 0.5, with the greatest discrepancy occurring for inner-shell analogs (high $\zeta$). High $\zeta$ values correspond to deeper initial potentials where the “quantum pressure” ($T_{\text{local}}\sim\zeta^{2}$) significantly skews the average energy relative to the experimental peak.
V Scaling Laws and the Energy Detection Discrepancy
A critical quantitative observation is found in the integrated energy values shown in the plot headers. In both cases, the peak energy $E_{\text{peak}}$ is significantly smaller than the total scattering energy $E_{\text{scatter}}$ (the integrated average $\langle E\rangle$). Specifically, for the $\zeta=1,Q=1$ case, the ”energy conservation” value $\langle E\rangle$ is driven higher by the potential energy density near the origin, whereas the experimental peak corresponds to the most probable radial position. This discrepancy validates our hypothesis that fragment detectors fundamentally probe the radial geometry of the initial state rather than the integrated energy mean.
@@ V Scaling Laws and the Energy Detection Discrepancy
V.1 Influence of Repulsive Charge $Q$
@@
V.2 Orbital Exponent $\zeta$: Inner-Shell vs. Outer-Shell Dynamics
@@
V.3 Universal Contour Map and Numerical Stability
@@
V.4 The Dimensionless Master Curve and Scaling Symmetry
The emergence of the 0.33 landmark is rooted in a fundamental scaling symmetry of the three-dimensional fragmentation dynamics. To characterize this, we examine the peak-to-mean ratio $R_{E}$ across the parametric landscape of mass ($M$), orbital localization ($\zeta$), and repulsive charge ($Q$). As demonstrated in Fig. 3, the ratio is governed by a single dimensionless coupling constant $\alpha$ when satisfying:
$\alpha=\frac{MQ}{\zeta}.$ (9)
Figure 3: Universal Scaling and the 0.33 Geometric Landmark. (Top) $R_{E}$ vs. $Q$ for the case $\zeta=Q,M=1,n=1$, showing a stable plateau at 0.33 (dashed line). (Middle) $R_{E}$ vs. $\zeta$ for the case $Q=1,M=\zeta,n=1$, exhibiting identical stabilization at 0.33. (Bottom) 3D contourf representation of the parameter space $(\zeta,Q)\in[1,5]$ with the mass scaling $M=\alpha\zeta/Q$. The emergence of three parallel, invariant planes for $\alpha=0.5$ ($R_{E}\approx 0.26$), $\alpha=1.0$ ($R_{E}\approx 0.33$), and $\alpha=2.0$ ($R_{E}\approx 0.39$) confirms that the energy detection efficiency is a pure function of the dimensionless scaling factor $\alpha$.
By substituting our proposed scaling relation (Eq. 9) into the kinetic term, we obtain:
$\langle T\rangle\propto\frac{\zeta^{2}}{\alpha\zeta/Q}=\frac{1}{\alpha}\zeta Q.$ (10)
Crucially, under this transformation, both the kinetic and potential energy components scale linearly with the product $\zeta Q$. This synchronization ensures that the relative weights of the energy components—and thus the geometric morphology of the resulting distribution P(E)P(E)—depend solely on the dimensionless coupling constant $\alpha$. This is also illustrated in the upper figure of Fig. 3, where we plot the $P(E)$ curve with scaled parameter $\frac{E}{Q\zeta}$. We found the base curve with variant $Q$ where $\zeta=M=n=1$, matches eactly with curves of variant $\alpha$ (the parameters satisfying Eq. 9). As can be illustated with the exact match of base curve $Q=1$ with $\alpha=1$ ($Q=2,\zeta=1$ and $Q=1,\zeta=2$), and base curve $Q=2$ with $\alpha=2$ ($Q=2,\zeta=1$ and $Q=1,\zeta=2$). The scaled shape of $P(E)$ is controlled by the parameter $\alpha$, where higher $\alpha$ values result in a narrower distribution with a constant peak position. This behavior provides an intuitive representation of the $\alpha$-dependent scaling ratio. By using this distribution as a baseline to fit experimental data, one can uncover fundamental properties like $M$, $\zeta$ and $Q$. We demonstrate this utility later by comparing our model to the beta decay spectrum.
One step further, this shows $R_{E}$ remains invariant along the $M\propto\zeta$ trajectories: the system undergoes a self-similar expansion in energy space where the $4\pi r^{2}$ masking effect is preserved as a constant geometric ratio. The invariance observed in the parallel planes of Fig. 3 (lower figure) can be derived by examining the homogeneous scaling properties of the system’s energy components. For a 1s Slater orbital, the expectation value of the potential energy scales linearly with the orbital exponent, $\langle V\rangle\propto Q\zeta$, while the expectation value of the kinetic energy (quantum pressure) scales quadratically with the exponent and inversely with the mass, $\langle T\rangle\propto\zeta^{2}/M$.
To determine the universal values of $R_{E}$ across the scaling parameter $\alpha$, we utilize the invariance property to map the ratio onto our previously computed $Q$-dependence results (where $M=\zeta=n=1$) from Fig. 2:
$R_{E}(\alpha)=R_{E}(Q)\rvert_{M=\zeta=1}.$ (11)
The figure is also re-ploted in Fig. 4,
Figure 4: $R_{E}$ vs. $\alpha$ for the case, same as computations with baseline for Q $M=\zeta=n=1$.
with numerical behavior of $R_{E}$ is thus characterized by the following limits:
$\lim_{\alpha\to 0}R_{E}=0,\quad\lim_{\alpha\to\infty}R_{E}=0.5.$ (12)
When $\alpha=1$, the system achieves a “geometric equilibrium” where the $4\pi r^{2}$ volume element sequesters exactly two-thirds of the total energy in the high-momentum radial tail, yielding the 0.33 landmark. As $\alpha$ decreases (physically representing high $Q$ or low $M$), the ratio monotonically climbs toward the 0.5 universal bound, a regime where the repulsive field dominates the initial-state quantum pressure.
Conversely, as $\alpha$ increases (high localization $\zeta$), $R_{E}$ decays precipitously, falling below 0.2 as $\zeta\to 5$ and asymptotically approaching 0. This signals a profound “detection crisis” where the vast majority of the interaction energy becomes experimentally invisible to peak-centric calorimetry. This confirms that for sufficiently localized states—such as core-shell electrons or compressed nucleons—the true interaction energy is effectively masked by the topological constraints of 3D space. Consequently, the mapping $R_{E}(\alpha)$ serves as a universal master curve for energy reconstruction across all scales of 3D quantum fragmentation.
VI Principal Quantum Number Effects and Multi-Peak Phenomena